OpenStudy (anonymous):

I am doing square roots.I have 3 x-20 square +9=36 how do I solve this symbolically?

OpenStudy (anonymous):

can you draw it.. ? :(

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

\[3\sqrt{x-20}+9=36\]

OpenStudy (anonymous):

\[i came up with 3\sqrt{x-7}\]

OpenStudy (anonymous):

@nelsonc12 that looks perfect ;)... \[\Large 3\sqrt{x-20}=36-9\] \[\Large 3\sqrt{x-20}=27\] now divide by 3and you'll have \[\Large \sqrt{x-20}=\frac{27}{3}\] \[\Large \sqrt{x-20}=9\] now square both sides ... \[\Large x-20=9^2 \]can you do it now ?

OpenStudy (anonymous):

ok... hold on ;)

OpenStudy (anonymous):

\[\LARGE x=9^2+20\] \[\LARGE x=81+20\] how about now ? :)

OpenStudy (anonymous):

That's correct. Well done ;)

OpenStudy (anonymous):

thank you do you know anything about f(x) to find domain and range

OpenStudy (anonymous):

\[f(x)=\sqrt{?81-x^2}\]

OpenStudy (anonymous):

post your questions on the wall not here... there should be someone to help if I'm not around...

OpenStudy (anonymous):

okay thank youj

OpenStudy (anonymous):

what do you have to find here?

OpenStudy (anonymous):

\[\LARGE f(x)=\sqrt{81-x^2}\] do you mean this ?

OpenStudy (anonymous):

hmm.. what is range now !! O_O ... let me go to google translate lol :$

OpenStudy (anonymous):

well I'm not sure if I'm capable to do it, (I might be wrong something...) you better ask your question on the wall... although I think \[\LARGE f(x)=\sqrt{81-x^2}\] \[\LARGE x\in [-9,9]\] anyway to get sure post it on wall... Good luck !